### v1

# On taming Moffatt-Kimura vortices of doom in the viscous case [^1] Dedicated to Professor Peter Constantin on the occasion of his 75th birthday

Abstract
In this note we propose a two-layer viscous mechanism for preventing finite time singularity formation in the Moffatt-Kimura model of two counter-rotating vortex rings colliding at a nontrivial angle. In the first layer the scenario is recast within the framework of the study of turbulent dissipation based on a suitably defined ‘scale of sparseness’ of the regions of intense fluid activity. Here it is found that the problem is (at worst) critical, i.e., the upper bound on the scale of sparseness of the vorticity super-level sets is comparable to the lower bound on the radius of spatial analyticity. In the second layer, an additional more subtle mechanism is identified, potentially capable of driving the scale of sparseness into the dissipation range and preventing the formation of a singularity. The mechanism originates in certain analytic cancellation properties of the vortex-stretching term in the sense of compensated compactness in Hardy spaces which then convert information on local mean oscillations of the vorticity direction (boundedness in certain log-composite weighted local bmo spaces) into log-composite faster decay of the vorticity super-level sets.
1 Introduction
In the study of possible singularity formation in the 3D Euler and Navier-Stokes (NS) equations the scenarios based on setting up a configuration of vortex structures (e.g., vortex tubes or vortex rings) colliding at a nontrivial angle and engineered with an eye on maximizing the vorticity amplification primarily via the mechanism of vortex stretching have been of a particular interest as plausible avenues to arrive at a singularity.
Moffatt and Kimura [15, 16] proposed a model of two counter-rotating (opposing circulation) vortex rings colliding at a nontrivial angle for which they derived a system of ODEs relating the key physical quantities of interest. The mechanism is essentially Eulerian and the effects of the viscosity are discussed a posteriori. Dynamics is divided in two phases, Phase I takes place before the reconnection and Phase II during reconnection.
Figure 1: initial configuration [17]
While the scenario remains a plausible finite time blow-up scenario in the inviscid case, the situation in the viscous case, due to the process of viscous reconnection, is less transparent. Yao and Hussain [20] performed a DNS study initialized at the Moffatt-Kimura initial configuration and observed that – in Phase II – there seems to be significant flattening and stripping of the vortex cores which would be inconsistent with the model. Moreover they argued that the formation of bridges during the viscous reconnection is capable of arresting the growth of the vorticity, a mechanism that was not accounted for in the model. Moffatt and Kimura addressed some of these issues in [17] and – in particular – pointed out that according to their updated calculations the conditions needed to attain a significant amplification of the vorticity magnitude are far beyond what can be realized either in experiments or numerical simulations.
Figure 2: viscous reconnection, bridges and threads at $\emph{Re}=4000$ [20]
The goal of this note is to consider the problem from a slightly different angle and propose a two-layer mechanism for preventing a finite time blow-up in the viscous case.
The underlying mechanism is the one of turbulent dissipation considered in the mathematical framework based on the suitably defined ‘scale of sparseness’ of the vorticity super-level sets quantifying spatial intermittency. Shortly, if the scale of sparseness near a possible singular time falls below the scale of the radius of spatial analyticity measured in $L^{\infty}$ (a natural dissipation scale quantifying analytic smoothing present in the viscous case), the harmonic measure maximum principle will prevent a further growth of the vorticity magnitude contradicting a finite time blow-up [12, 3, 13, 14].
In the first layer we observe that Moffatt-Kimura scenario is (at worst) critical within the aforementioned framework, i.e., the scale of sparseness encoded in the model (diameter of the vortex cores as they rush toward the ‘tipping point’) and the scale of the radius of spatial analyticity coincide. This is based on the assumption made in the original model where the vortex cores remain mostly compact and nearly circular even when transitioning to Phase II. Note that within the framework, this is the geometrically worst case scenario since flattening and stripping of the cores – as observed in the DNS performed by Yao and Hussain – could possibly drive the scale of sparseness into the dissipation range.
Since any quantification of the possible gain of sparseness due to the formation of bridges and threads in the process of viscous reconnection is out of reach, we propose to explore a different route to (possibly) braking the criticality – the one based on the study of local mean oscillations of the vorticity direction.
The significance of the vorticity direction in the study of the problem of global regularity for the 3D NS system was first recognized by Constantin in [6] where a singular integral representation of the stretching factor in the evolution of the vorticity magnitude was derived. The representation features a geometric kernel the strength of which is depleted precisely by local coherence of the vorticity direction. In a follow-up work [7] Constantin and Fefferman showed that as long as the vorticity direction is Lipschitz in the regions of intense vorticity no finite time blow-up can occur. Subsequently Beirao Da Veiga and Berselli reduced the Lipschitz condition to $\frac{1}{2}$-Hölder [1] which was – in turn – followed up by several works including a spatiotemporal localization of the $\frac{1}{2}$-Hölder regularity criterion [11] and a work on the role of local coherence of the vorticity direction in the study of 3D enstrophy cascade in the NS flows [8].
In a particular case of the critical blow-up, a significant reduction of regularity of the vorticity direction needed to prevent a finite time blow-up is possible. More precisely Giga and Miura in [10] demonstrated that in the case of Type I blow up (a blow-up rate of the $L^{\infty}$-norm is bounded by the self-similar rate) it suffices that the vorticity direction is uniformly continuous in the spatial variable.
Since the spatiotemporal scaling of Moffatt-Kimura scenario is consistent with a Type I blow-up, a potential singularity formation at the tipping point would then force the regularity of the vorticity direction below that of uniform continuity. In search of a suitable functional class it is helpful to notice that dynamics of the viscous reconnection at high Reynolds numbers seems to feature a creation of secondary vortices (an ‘avalanche’) which then tend to tangle up in intricate ways (Figure 3). Consequently – as the flow approaches a possible singularity – the vorticity direction is expected to exhibit a nontrivial oscillatory behavior and the class of local weighted spaces of functions of bounded mean oscillations seems a reasonable choice to encode this.
Figure 3: viscous reconnection, avalanche at $\emph{Re}=40000$ [21]
Let us denote the mean oscillation of a function $f$ over the cube $I=I(x,r)$ centered at $x$ with the side length $r$ by
$\Omega(f,I(x,r))=\frac{1}{|I(x,r)|}\int_{I(x,r)}|f(x)-f_{I}|\,dx$
where $f_{I}$ is the mean value of $f$ over $I$. Assuming that $f\in L^{1}$ we can focus on small scales, say $0<r<\frac{1}{2}$, and define the local weighted space of functions of bounded mean oscillations $\widetilde{bmo}_{\phi}$ as follows. For a positive non-decreasing function $\phi$ on $(0,\frac{1}{2})$ require that the quantity
$\|f\|_{\widetilde{bmo}_{\phi}}=\|f\|_{L^{1}}+\sup_{x\in\mathbb{R}^{3},0<r,\frac{1}{2}}\frac{\Omega(f,I(x,r))}{\phi(r)}$
is finite. In the case of a power weight $\phi(r)=r^{\alpha}$ for $0<\alpha\leq 1$ it is known that being in $\widetilde{bmo}_{\phi}$ is equivalent to being in the local $\alpha$-Hölder class.
Since we are already in a critical scenario, the weights of the most interest here are the ones that would allow for discontinuities (due to Giga and Miura). One can show that $\widetilde{bmo}_{\phi}$ contains discontinuous functions if and only if
$\int_{0}^{\frac{1}{2}}\frac{\phi(r)}{r}\,dr=\infty$
and since the vorticity direction (being bounded) is trivially in $\widetilde{bmo}_{1}$ the sequence of weights
$\phi_{k}(r)=\frac{1}{\log^{k}(|\log r|)}$
($\log^{l}$ denotes an l-fold composition) generates a reasonable scale of spaces to focus on (there are functions that grow at infinity slower than any log-composite and are for all practical purposes equivalent to a constant, e.g, inverse Ackerman function; for the time being it seems sensible to defocus from this).
This setting also points to another, perhaps less obvious contrast between inviscid and viscous Moffatt-Kimura scenarios (compared to no smoothing vs. analytic smoothing). In the inviscid case – due to the lack of the shear stress – the scenario of two counter-rotating vortex rings rushing to the tipping point at a nontrivial angle is consistent with a formation of a simple jump discontinuity in the vorticity direction. Note that this is precisely a representative of the case where a function would be in $\widetilde{bmo}_{1}$ (in this case a trivial bound) but not in any of $\widetilde{bmo}_{\phi}$ with the log-composite weights (in one dimension, think about the signum function on (r,r)(-r,r), the mean oscillation is exactly one, no dependence on $r$). In contrast, in the viscous case, it is not unreasonable to expect that the shear stress-driven component of spatial dynamics near the potential singularity – grinding of topologically intricate local fluid layers (Figure 3) – would be capable of taming the local mean oscillations just enough to nudge them over the trivial bound and into the log-composite $\widetilde{bmo}_{\phi}$ realm.
In the second layer we observe that – in the viscous Moffatt-Kimura scenario – previous work [4, 9] based on utilizing analytic cancellations in the vortex-stretching term in the sense of compensated compactness in Hardy spaces to (logarithmically) improve upon the a priori bound on the $L^{1}$-norm of the vorticity obtained by Constantin [5] could be used to (logarithmically) improve upon the scale of sparseness, breaking the criticality noted in the first layer. The key assumption in [9] was precisely that the vorticity direction belonged to one of the log-composite $\widetilde{bmo}_{\phi}$ spaces.
The note is organized as follows. Section 2 is devoted to the criticality scenario (layer one), Section 3 to depicting a possible road to the logarithmic sub-criticality (layer two), and Section 4 presents a rigorous dissipation argument.
2 Criticality
On one hand, recall that in Moffatt-Kimura scenario of two counter-rotating vortex rings of the constant radii $R$ colliding at a nontrivial angle the maximal vorticity amplification rate at the tipping point (potential singularity) is given by
$\mathcal{A}_{\omega}=\frac{\omega_{max}(t_{c})}{\omega_{0}}=\frac{\delta^{2}_{0}}{\delta^{2}(t_{c})}$
where $\delta_{0}$ is the diameter of the vortex cores in the initial configuration, $\omega_{0}$ the initial axial velocity given by ${\omega_{0}=\frac{\Gamma}{4\pi\delta^{2}_{0}}}$ ($\Gamma$ and $\Gamma$ are the initial circulations), $\delta(t_{c})$ the diameter of the vortex cores at the critical time (tipping point), and $\omega_{max}(t_{c})$ the vorticity maximum at the critical time [17]. Solving for $\delta(t_{c})$ gives
$\delta(t_{c})=\biggl(\frac{\Gamma}{4\pi}\biggr)^{\frac{1}{2}}\frac{1}{\|\omega(t_{c})\|_{\infty}^{\frac{1}{2}}}.$
$\delta(t)$ is a natural small (spatial) scale intrinsic to the model, the macro scale being the radii of the rings $R$. It is also comparable to the scale of sparseness of the vorticity super-level sets.
On the other hand, recall that a lower bound on the radius of spatial analyticity of the vorticity field is given by [3]
$\rho(t)=\frac{1}{c_{*}}\frac{\nu^{\frac{1}{2}}}{\|\omega(t)\|_{\infty}^{\frac{1}{2}}}$
where $\nu$ is the viscosity.
In order to prevent the blow-up in this framework one needs ${\rho(t)\geq\delta(t)}$ as $t$ approaches $t_{c}$ and since the dynamic quantities coincide (we are in a critical scenario) the scale of sparseness will fall into the dissipation range if
$\frac{\Gamma}{\nu}\leq\frac{4\pi}{c_{*}^{2}},$
i.e., only when the Reynolds number of the initial configuration is of order 1 (the battle of constants is not going all that great). However, since the scale of sparseness noted here is based on a stipulation that the cores remain mostly compact and nearly circular, an assumption made in [16] but also challenged in [20] due to flattening and stripping of the cores during the viscous reconnection, there may be room for a dynamic improvement.
3 Beyond criticality?
Let us start by observing that the scale of sparseness presented in the previous section can alternatively be obtained from the basic geometry of the model and an a priori bound on the decay of the volume of the vorticity super-level sets derived from the vorticity-velocity formulation of the NS system.
Recall that the vorticity analogue of Leray’s a priori $L^{2}$ bound on the velocity is Constantin’s a priori $L^{1}$ bound derived in [5]. The $L^{1}$ bound on the vorticity combined with a trivial bound on the distribution function,
$|\{x\in\mathbb{R}^{3}:\,|\omega(x,t)|>M\}|\leq\frac{\|\omega(t)\|_{1}}{M}$
yields the following bound on the volume of the vorticity super-level sets relevant in the framework ($\lambda$ is a positive constant less than one the choice of which depends on some other tuning parameters; a complete argument of this type will be given in Section 4)
$|\{x\in\mathbb{R}^{3}:\,|\omega(x,t)|>\lambda\|\omega(t)\|_{\infty}\}|\leq\frac{c(\lambda,\|\omega_{0}\|_{1},T)}{\|\omega(t)\|_{\infty}}$
for any $t$ in an interval $(0,T)$. Since the radii of the rings $R$ represent the macro scale in the model, the decay of the volume given by the above inequality as the flow approaches a possible singularity will force the decay of the diameters of the vortex cores comparable to ${\frac{1}{\|\omega(t)\|_{\infty}^{\frac{1}{2}}}}$ – the same small scale that appeared in the previous section (as before this is – from the point of view of sparseness – geometrically worst case scenario, the observed flattening and stripping of the cores would only make it sparser).
It is this alternative way of deriving an upper bound on the diameters of the vortex cores that is amenable to further analysis – more precisely, a logarithmic improvement of the $L^{1}$ bound on the vorticity would imply a logarithmic improvement of the rate of decay of the volume of the vorticity super-level sets which would – in turn – yield a logarithmic improvement on the upper bound on the vortex core diameters, breaking the criticality. In other words, the problem is now reduced to obtaining a logarithmic improvement of the $L^{1}$ bound on the vorticity, a question of its independent interest.
The first contribution in this direction was presented in [4] where it was shown that as long as the vorticity direction is in ${\widetilde{bmo}_{\frac{1}{|\log r|}}}$ the vorticity magnitude will remain in $LLogL$. The proof was based on the duality between local Hardy and BMO spaces, analytic cancellations in the vortex-stretching term in the sense of the ‘div-curl’ lemma in Hardy spaces, some results on multipliers in local BMO spaces and Coifman-Rochberg $BMO$ estimate on the logarithm of the maximal function.
A natural follow up question was whether it was possible to generalize this result to the log-composite weights
$\phi_{k}(r)=\frac{1}{\log^{k}(|\log r|)},$
ideally for any positive integer $k$. Unfortunately there is an obstruction – there is no analogue of Coifman-Rochberg estimate in the weighted $BMO$ spaces (in the case of log-composite weights one can construct several types of counterexamples). However, it is possible to derive a ‘dynamic version’ of Coifman-Rochberg for a family of time-dependent functions bounded above and below by algebraic rates, modeling a flow approaching a possible singularity (with no restrictions on the strength of the singularity) [9]. This was then used to obtain a generalization of the $LLogL$ result – it was demonstrated that as long as the vorticity direction belongs to ${\widetilde{bmo}_{\phi_{k}(r)}}$ and the blow-up is of an algebraic type, the vorticity magnitude will stay in $L\frac{1}{\phi_{k}(L)}$. Similarly as in the $L^{1}$ case, this implies the following bound on the decay of the volume of the vorticity super-level sets of interest,
$|\{x\in\mathbb{R}^{3}:\,|\omega(x,t)|>\lambda\|\omega(t)\|_{\infty}\}|\leq c(\lambda,\|\omega_{0}\|_{L\frac{1}{\phi_{k}(L)}},T)\,\frac{\phi_{k}(\|\omega(t)\|_{\infty})}{\|\omega(t)\|_{\infty}}$ (3.1)
for any $t$ in an interval $(0,T)$.
In summary, as long as the local mean oscillations of the vorticity direction indeed exhibit a log-composite decay near the potential singularity the flow remains in a sub-critical regime and a finite time blow-up is prevented.
4 Dissipation via the harmonic measure maximum principle
The purpose of this section is to provide a rigorous argument showing that as long as the bound (3.1) holds, the $L^{\infty}$ norm of the vorticity will remain bounded and no singularity will form.
Let us start with recalling definitions of local sparseness at scale suitable for the mathematical analysis of spatial intermittency [12].
Definition 4.1.
For a spatial point $x_{0}$ and $\delta\in(0,1)$, an open set $S$ is 1D $\delta$-sparse around $x_{0}$ at scale $r$ if there exists a unit vector $\nu$ such that
$\frac{|S\cap(x_{0}-r\nu,x_{0}+r\nu|}{2r}\leq\delta\ .$
Definition 4.2.
For a spatial point $x_{0}$ and $\delta\in(0,1)$, an open set $S$ is 3D $\delta$-sparse around $x_{0}$ at scale $r$ if
$\frac{|S\cap B_{r}(x_{0})|}{|B_{r}(x_{0})|}\leq\delta\ .$
Note that local 3D $\delta$-sparseness at a scale implies 1D $(\delta)^{\frac{1}{3}}$-sparseness at the same scale; the converse is false.
The sets of interest will be the vorticity super-level sets – more precisely – we will consider the following. For a time $t$ and a spatial point $y\in\mathbb{R}^{3}$ there exists a vorticity component $\omega^{i,\pm}$ such that $\omega^{i,\pm}(y,t)=|\omega(y,t)|$ (opting for the max norm in $\mathbb{R}^{3}$). The associated super-level set is then defined by
$V^{i,\pm}_{t}(y)=\{x\in\mathbb{R}^{3}:\,\omega^{i,\pm}(x,t)>\lambda\|\omega(t)\|_{\infty}\}$ (4.1)
where $\lambda\in(0,1)$ will be a suitably chosen parameter.
Local sparseness at scale will be utilized via the harmonic measure maximum principle. Due to the rotational invariance of the equations and locality of the argument one can always assume that the direction of local 1D sparseness is a coordinate direction. Hence the setting of one complex variable suffices (in several complex variables, the harmonic measure would be replaced with the plurisubharmonic measure).
Proposition 4.3 ([18]).
Let $\Omega$ be an open, connected set in $\mathbb{C}$ such that its boundary has nonzero Hausdorff dimension, and let $K$ be a Borel subset of the boundary. Suppose that $u$ is a subharmonic function on $\Omega$ satisfying
$u(z)$ $\leq M\ ,\quad\textrm{for }z\in\Omega$
$\limsup_{z\to\zeta}u(z)$ $\leq m\ ,\quad\textrm{for }\zeta\in K.$
Then
$u(z)\leq m\,h(z,\Omega,K)+M(1-h(z,\Omega,K))\ ,\quad\textrm{for }z\in\Omega.$
(Here, $h(z,\Omega,K)$ denotes the harmonic measure of $K$ with respect to $\Omega$, evaluated at $z$.)
The following extremal property of the harmonic measure in the unit disc $\mathbb{D}$ will be helpful in the calculation.
Proposition 4.4 ([19]).
Let $\alpha$ be in $(0,1)$, $K$ a closed subset of [1,1][-1,1] such that $|K|=2\alpha$, and suppose that the origin is in $\mathbb{D}\setminus K$. Then
$h(0,\mathbb{D},K)\geq h(0,\mathbb{D},K_{\alpha})=\frac{2}{\pi}\arcsin\frac{1-(1-\alpha)^{2}}{1+(1-\alpha)^{2}}$
where $K_{\alpha}=[-1,-1+\alpha]\cup[1-\alpha,1]$.
Lastly, we need a local-in-time lower bound on the radius of spatial analyticity of the vorticity field in $L^{\infty}$.
Theorem 4.5 ([3]).
Let $\omega_{0}\in L^{\infty}$ and $M\in(1,\infty)$. Then there exists a unique mild solution to the vorticity formulation of the 3D HD NS system $\omega\in C_{w}([0,T],L^{\infty})$ where
$T\geq\frac{1}{c_{1}(M)}\frac{\nu}{\|\omega_{0}\|_{\infty}}$
and for any $t\in(0,T]$ the solution $\omega$ is the $\mathbb{R}^{3}$-restriction of a holomorphic function $\omega$ defined in
$\Omega_{t}=\biggl\{x+iy\in\mathbb{C}^{3}:\,|y|<\frac{1}{c_{2}(M)}t^{\frac{1}{2}}\biggr\}$
satisfying $\|\omega(t)\|_{L^{\infty}(\Omega_{t})}\leq M\|\omega_{0}\|_{\infty}$.
In what follows the choice of the cut-off parameter $\lambda$ and the sparseness parameter $\delta$ will have to be consistent with the choice to be made in the application of the harmonic measure maximum principle – with that in mind, set $\delta=\frac{3}{4}$ and $\lambda=\frac{1}{2M}$ where $M$ is the solution to
$\frac{1}{2}h^{*}+(1-h^{*})M=1\ \ \mbox{and}\ \ h^{*}=\frac{2}{\pi}\arcsin\frac{1-\frac{3}{4}^{\frac{2}{3}}}{1+\frac{3}{4}^{\frac{2}{3}}}$ (4.2)
and – at the same time – $M$ to be used in the application of Theorem 4.5 (it is easy to check that $M>1$).
In addition, it will be convenient to have a designation of ‘escape time’. Let $\omega\in C_{w}\bigl([0,T^{*}),L^{\infty}\bigr)$ where $T^{*}$ is the first blow-up time. A time $t$ is an escape time provided $\|\omega(s)\|_{\infty}>\|\omega(t)\|_{\infty}$ for any $s\in(t,T^{*})$ (local-in-time well-posedness in $L^{\infty}$ implies that for any level there exists a unique escape time).
All tools for the proof of the main result are now collected.
Theorem 4.6.
Consider Moffatt-Kimura scenario in the viscous case and suppose that the a priori bound (3.1) holds. Then the vorticity magnitude remains bounded and a finite time blow-up is avoided.
Proof.
Let $M$ be as in (4.2) and $t$ an escape time. Solve the 3D NS system at $t$ according to Theorem 4.5 with the same choice of $M$, and let $s=t+T_{t}$ where $T_{t}$ is the maximal time of existence guaranteed by the theorem. The the solution at $s$ is analytic with the uniform radius of analyticity of at least
$\frac{1}{c_{3}}\frac{\nu^{\frac{1}{2}}}{\|\omega(t)\|_{\infty}^{\frac{1}{2}}};$
since $t$ is an escape time the lower bound at $s$ could be replaced with
$\rho_{s}=\frac{1}{c_{3}}\frac{\nu^{\frac{1}{2}}}{\|\omega(s)\|_{\infty}^{\frac{1}{2}}}.$ (4.3)
Next, note that the bound (3.1) (taking any $T$ greater than the potential blow-up time) and the geometry of Moffatt-Kimura scenario imply that for any $y\in\mathbb{R}^{3}$ the scale of 3D sparseness of the super-level sets $V^{i,\pm}_{s}(y)$ defined in (4.1) is given by
$r_{s}=c_{4}(\|\omega_{0}\|_{L\frac{1}{\phi_{k}(L)}},R)\,\biggl(\frac{\phi_{k}(\|\omega(s)\|_{\infty})}{\|\omega(s)\|_{\infty}}\biggr)^{\frac{1}{2}}$
(with the choice of the parameters made in (4.2).
At this point we make a choice of the escape time $t$ to be an escape time for which $\rho_{s}\geq r_{s}$; then $\omega(s)$ is both analytic and 1D $\frac{3}{4}^{\frac{1}{3}}$-sparse at scale $r_{s}$.
Let $x_{0}\in\mathbb{R}^{3}$ be arbitrary. We aim to show $|\omega(x_{0},s)|\leq\|\omega(t)\|_{\infty}$; this would contradict $t$ being an escape time and conclude the argument.
Due to the translational and rotational invariance of the equations, we can assume that $x_{0}$ is the origin and the direction of local 1D sparseness is the coordinate direction $e_{1}$. Immerse $e_{1}$ in the complex plane and consider
$D_{r_{s}}=\{z\in\mathbb{C}:\,|z|<r_{s}\}.$
Since each $\omega_{i}^{\pm}(\cdot,s)$ is subharmonic on $D_{r_{s}}$ the stage is set for an application of the harmonic measure maximum principle (Proposition 4.3). Let $\omega_{i}^{\pm}$ be the local (at 0) maximal component, i.e., $\omega_{i}^{\pm}(0,s)=|\omega(0,s)|$, and recall that the corresponding super-level set $V_{s}^{i,\pm}=V_{s}^{i,\pm}(0)$ is 1D $\frac{3}{4}^{\frac{1}{3}}$-sparse at scale $r_{s}$.
Next, define a compact set $K$ to be the complement in $r_{s},r_{s}]$ of the set $V_{s}^{i,\pm}\cap(-r_{s},r_{s})$, and note that – due to sparseness – $|K|\geq 2r_{s}\bigl(1-\frac{3}{4}^{\frac{1}{3}}\bigr)$.
In order to apply the estimate on the harmonic measure given in Proposition 4.4, we need to consider the case $0\in K$ separately. This is straightforward since in this case
$|\omega(0,s)|=\omega_{i}^{\pm}(0,s)\leq\frac{1}{2M}\|\omega(s)\|_{\infty}\leq\frac{1}{2}\|\omega(t)\|_{\infty}$
and we obtain a contradiction with $t$ being an escape time (the last inequality follows from the bound on the vorticity given in Theorem 4.5).
In the case $0\notin K$, since the harmonic measure is invariant with respect to $z\to\frac{1}{r_{s}}z$, Proposition 4.4 yields
$h(0,D_{r_{s}},K)\geq\frac{2}{\pi}\arcsin\frac{1-\frac{3}{4}^{\frac{2}{3}}}{1+\frac{3}{4}^{\frac{2}{3}}}$
which is precisely $h^{*}$ in (4.2). Hence, Proposition 4.3 implies the following bound
$|\omega(0,s)|=\omega_{i}^{\pm}(0,s)\leq h^{*}\frac{1}{2}\|\omega(t)\|_{\infty}+(1-h^{*})M\|\omega(t)\|_{\infty}=\|\omega(t)\|_{\infty}$
by the choice of parameters made in (4.2), and we obtain a contradiction again. This completes the proof.
5 Conclusion
The main goal of this note was to propose a viscous two-layer mechanism for avoiding a finite time singularity formation in Moffatt-Kimura scenario of two counter-rotating vortex rings approaching a tipping point at a non-trivial angle. It is worth emphasizing that being in the viscous case is crucial for both layers.
The first layer is built on having a lower bound on the radius of spatial analyticity stemming from local-in-time analytic smoothing of the NS system. In the Euler case there is no smoothing and the only way to generate a local-in-time analytic solution is to start with the analytic initial data but even then the lower bound is inadequate (decreasing instead of increasing).
The second layer hinges on a stipulation that the shear stress component of topologically intricate dynamics of the viscous reconnection – especially at high Reynolds numbers (Figure 3) – is capable of slowing down the local oscillations of the vorticity direction just enough to tip them over from being bounded in mean (a trivial bound) to having a log-composite decay in mean where the number of composites can be arbitrary large. This is in contrast with the inviscid case where – due to the lack of the shear stress – the vorticity direction in Moffatt-Kimura scenario is expected to exhibit a simple jump discontinuity and no decay of the local mean oscillations (saturating the trivial bound).
One last remark is to note that if this mechanism was realized in a particular flow, it might be quite impossible to detect in a computational simulation. Namely, if the number of log-composites was $k$, then the level of the vorticity magnitude at which the crossover into the sub-critical regime took place would be of the order of a tetration of height $k$ (with base $e$ and the Reynolds number on top), and the spatial scale to which the flow would have to be resolved would be of the order of its reciprocal. In other words, a simulation run would likely indicate that the flow remained in the critical regime in which the battle of constants was already lost pointing to a finite time blow-up.
6 Acknowledgments
The work is supported in part by the National Science Foundation grant DMS 2307657. The author would like to express his gratitude to Professor Peter Constantin for opening the door (quite emphatically) into the realm of the mathematical study of the interplay between the geometric properties of 3D incompressible flows and their regularity, as well as for his support over the years.